A Complete Bibliography of ACM Transactions on Spatial Algorithms and Systems (TSAS)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

02 March 2022 Version 1.06

Title word cross-reference

2 [ABM16, WK18]. 3 [ABM16, VTSD18]. A^{⋆} [PWH21]. K [PAB^{+}16, CCBS18].

2017 [Are19]. 2019 [Are21].

Accessible [KKT^{+}18]. Accessing [CSF^{+}19]. accident [HHRR20]. Accumulated [KS15]. Accurate [ABY17, BBS19, WLL^{+}19]. ACM [Are19, Are21]. Activity [MH19]. Activity-aware [MH19]. Adaptive [DAGM21, DPG20]. Adversarial [WCC^{+}20]. Against [GDSB16]. Agents [KKT^{+}18]. Aggregate [DCY^{+}18]. Aggregation [PWH21, VTSD18]. AI [Wer21]. AI-Enhanced [Wer21]. Air [SKZ^{+}20]. Aircraft [ACS19]. Algorithm [BMNP16, BVW16, FAMF16, PLHC19]. Algorithms [GHN15, Gol19, KTY^{+}18, YKC20]. Along [MKW20]. among [EP20, GHPS21]. Analysis [BM20, LA19, MIF17, RLA19]. Analyzing [BPS18]. Angular [Pet21]. Anomalous [DPKW19]. Anomaly [SGBM20, WLL^{+}19]. Apache [WMPH19]. Application [JS19, TIKG18, TL20]. Applications [AWD^{+}18, FNCO20, SGBM20]. Applied [EP20]. Approach [ARF19, FNCO20, HBH^{+}21, KPS17, MKW20, TLF^{+}20]. Approaches [CDFP21]. Apps [CSKB19]. Area [BMVS16, PWH21, Pet21]. Area-Preserving [BMVS16].

Arrangements [PSD^{+}21]. Arterial [MKW20]. Assessment [LNK^{+}21]. Assigned [TSK15]. Assignment [SSTN19]. Assimilation [LH17]. Assisted [HKK^{+}19]. Attacks [GDSB16]. Attention [ZWT^{+}20]. Auto [STZ^{+}20, ZGP19]. Auto-Encoders [STZ^{+}20]. Auto-regressive [ZGP19]. Autologistic [SMM19]. Automatic [ZSBA21]. Aware [CKT^{+}19, GKR16, HYL16, DV21, MH19, NKTB20]. Axis [TIKG18].

Balancing [DAGM21]. Based [AFHW15, FNCO20, GDSB16, LH17, MAK^{+}18, MIF17, PK16, YÖR20, ZGP19, BM20, BASM21, MKW20, WCM20, YSWZ18]. Batch [CCBS18]. Behavior [HKK^{+}19, LZLL20, ZSFB20]. Behaviors [TKC17]. Best [Are19, Are21]. Big [DAGM21, VBME21]. Binary [TYZO15]. Boundary [GHN15]. Boundary-Labeling [GHN15]. Bucket [TYZO15]. Buy [AKRH19].

Candidates [NLC16]. Case [QKZU20]. Categorical [SMM19]. Cavity [CDFP21]. CellNet [MIF18]. Cellular [CSF^{+}19]. Centric [GDSB16]. Chains [PSD^{+}21]. Cities [ZSBA21]. City [DBB20]. Classification [CEGH17]. Closed [Pet21]. Closed-Form [Pet21]. Closure [PLHC19]. Cloud [BPM21, PBGA19]. Clustered [JS19]. Clustering [BASM21]. Co [AKAM17, PSD^{+}21, PAB^{+}16]. Co-occurrence [PAB^{+}16]. Co-Occurrences [AKAM17]. Collaborative [ATS^{+}16]. Collective [ISNU17]. Combination [TYZO15]. Combining [CDFP21]. Communities [FNCO20]. Comparison [AFHW15]. Complex [MF15]. Compression [PBGA19]. Computation [FAMF16, JS19, KKT^{+}18, KS15]. Computer [TLF^{+}20]. Computing [AKRH19, PSD^{+}21]. Conference [Are19, Are21]. Conflict [ACS19]. Congestion [CKT^{+}19]. Congestion-Aware [CKT^{+}19]. Conquer [MICNC21]. Consensus [ATS^{+}16]. Considering [ROOF17, YÖR20]. Consistent [CKM^{+}21]. Constraints [ZGP19]. Context [DV21, NKTB20, DCAA21]. Context-aware [DV21, NKTB20]. Continuous [ZYW^{+}21]. Contrast [LZLL20]. Control [ARF19]. Controlled [LZLL20]. Convolutional [HHRR20]. Coordinate [BDW18]. Coordinate-Free [BDW18]. Corridors [KTHK19]. Cost [KS15]. Coupled [SGBM20]. Crime [BPS18]. Crowd [CSF^{+}19, HBH^{+}21, HKK^{+}19, WLL^{+}19, WCC^{+}20]. Crowd-sourced [HBH^{+}21]. Crowds [NLC16]. Crowdsourced [BPS18, LNK^{+}21, SPKS16]. Crowdsourcing [CN20, LCKQ20, TSK15]. Curves [EP20]. Cusps [Pet21].

D [ABM16, VTSD18, WK18]. Data [AKM21, AH17, ACS19, BPS18, BDKS19, CSF^{+}19, CEGH17, DAGM21, DHQ^{+}21, DPG20, DV21, HBH^{+}21, ISNU17, KPS17, LH17, MR21, NKTB20, PBGA19, PLHC19, SGBM20, VTSD18, VBME21, Wer21, ZSBA21]. Data-driven [ACS19, DV21]. Database [SS21]. Databases [TL20]. Datasets [BMNP16, JSL^{+}20, PAB^{+}16, TDPT20, Wer21]. Decomposition [CDFP21]. Deep [DPG20, FNCO20, SKZ^{+}20, TL20, VBME21, WCM20, YKC20, ZSBA21]. Demand [ARF19]. Density [TLF^{+}20]. Dependency [ZGP19]. Deployment [DBB20]. Detail [VTSD18]. Detecting [FHK^{+}18]. Detection [ACS19, DPKW19, HHRR20, PLHC19, SGBM20, SOdB^{+}20, WLL^{+}19]. Determination [ROOF17]. Deviations [FHK^{+}18]. Diagrams [ZYW^{+}21].

Dictionary [PBGA19]. Differentiating [ZSFB20]. Direction [LA19]. Disaster [DBB20]. Discovery [NLC16]. Discrete [EP20]. Disk [MAK^{+}18]. Disk-Based [MAK^{+}18]. Distance [AFHW15]. Distances [EP20]. Distortion [Pet21]. Distributed [ARF19, DAGM21, TDPT20, WMPH19]. Distribution [BPS18]. Diversity [DHQ^{+}21]. Divide [MICNC21]. Drift [SOdB^{+}20]. driven [ACS19, DV21]. Driving [LZLL20, WLL^{+}19]. DropBox [DBB20]. Dynamic [DV21, EP20, GKR16]. Dynamics [QKZU20].

Efficient [ABY17, BBS19, DBB20, EEP16, FAMF16, KKT^{+}18, RHJC19, ROOF17, SS21, TYZO15]. Encoders [STZ^{+}20]. End [ZWT^{+}20]. End-to-End [ZWT^{+}20]. Energy [ABY17]. Energy-Efficient [ABY17]. Engulfs [BDW18]. Enhanced [KTY^{+}18, Wer21]. Enrichment [TL20]. Entities [SS21]. Environment [ATS^{+}16, TIKG18]. Environmental [FNCO20]. Epidemic [WCM20]. Equal [Pet21]. Equal-Area [Pet21]. Estimating [ISNU17, SKZ^{+}20]. Estimation [SPKS16, TLF^{+}20]. Euclidean [EP20]. Event [GDSB16, PK16, ZCLR16]. Event-Based [PK16]. Example [PSTT16]. Exchange [DBB20]. Experiment [LZLL20]. Expert [NLC16]. Exploring [SKZ^{+}20]. Extended [PAB^{+}16]. External [FAMF16, GDSB16].

Family [FDMW20]. Fast [JS19, LCKQ20]. Feature [YSWZ18]. Feature-based [YSWZ18]. Features [ZSFB20]. Fencing [TYZO15]. Filtering [LH17]. Finding [KTHK19, PWH21]. Fingerprint [LNK^{+}21]. Flash [TPZ15]. Flexible [MH19]. Flood [LA19, RLA19]. Flood-Risk [LA19]. Flow [ISNU17, WCC^{+}20]. Football [CEGH17]. Force [YÖR20]. Forecasting [WCM20, ZCLR16]. Form [Pet21]. Fragile [WK18]. Framework [AH17, ACS19, LCKQ20, MICNC21, TSK15]. Fréchet [EP20, GHPS21]. Free [BDW18].

GAN [WCC^{+}20]. GANs

[JSL^{+}20, SGBM20]. General [AH17]. Generalized [RHJC19]. Generating [JSL^{+}20, MF15]. Generation [BBS19, KPS17]. Generative [WCC^{+}20]. Generic [TPZ15]. Geo [AKM21, NLC16, TYZO15]. Geo-Fencing [TYZO15]. Geo-Located [NLC16]. Geo-Social [AKM21]. Geographical [FNCO20]. Geographically [BBS19]. Geography [NT18]. Geometric [BMNP16]. Geosensor [BDW18]. Geospatial [ZSFB20]. Global [Wer21, ZSBA21]. Global-scale [ZSBA21]. GloBiMapsAI [Wer21]. GPS [ABY17, FHK^{+}18, HBH^{+}21, MIF17, MIF18, SSTN19, YSWZ18]. GPS-Less [ABY17]. GPU [ABM16]. Graph [KTHK19, SS21]. Graphical [ISNU17]. Graphics [KS15]. Graphs [DPKW19]. Grid [MIF17]. Grid-Based [MIF17]. Grids [ABM16]. Group [KKT^{+}18, MH19, PK16, RHJC19]. GTS [LCKQ20]. guided [WCM20].

Hailing [JSL^{+}20]. Hashing [TYZO15]. Hierarchical [FDMW20, VTSD18]. Highway [TL20]. Historical [BVW16]. Human [DHQ^{+}21, HKK^{+}19, SGBM20, DCAA21, WLL^{+}19]. Hybrid [TYZO15].

Identifying [BDW18, FNCO20]. IGMM [SGBM20]. IGMM-GANs [SGBM20]. Images [GHN15, SKZ^{+}20]. Imbalance [ARF19]. Impact [CSKB19, DCAA21]. Improving [DBB20]. In-Bucket [TYZO15]. In-Route [CN20]. Incorporating [STZ^{+}20]. Incremental [DPG20]. Index [GHPS21, SS21, TPZ15]. Indexes [FDMW20]. Indexing [KTY^{+}18, MAK^{+}18].

Individual [SSTN19, DCAA21]. Indoor [LNK^{+}21, TKC17]. Inference [HBH^{+}21, ZSBA21]. Inferring [MIF18]. Influence [HYL16]. Influence-Aware [HYL16]. Information [AKRH19, DBB20, ROOF17, WCM20]. Infrastructure [BPM21]. Integer [PWH21]. Intelligent [HHRR20]. Intended [FHK^{+}18]. Interaction [YÖR20]. Interactive [BVW16]. Interchange [TL20]. Internet [NT18]. Interpolation [ABM16, MR21]. Interpretable [ZGP19]. Intersection [HHRR20]. Introduction [Are19, Are21, Gol19, YKC20]. Issue [Are19, Gol19, YKC20].

Join [TDPT20, WMPH19].

Keyword [AKM21].

Labeling [GHN15]. Labels [BVW16]. Large [LCKQ20, MICNC21]. Large-Scale [LCKQ20]. Layered [KPS17, TIKG18]. Learning [FNCO20, SKZ^{+}20, TL20, VBME21, WCM20, YKC20, ZGP19]. Learning-based [WCM20]. Less [ABY17]. Level [VTSD18]. Level-of-Detail [VTSD18]. LifeSteps [PSTT16]. Linear [PWH21]. Load [DAGM21]. Local [NLC16]. Localization [ABY17]. Located [NLC16]. Location [DHQ^{+}21, GDSB16, KKT^{+}18, PK16, SPKS16, WK18]. Location- [PK16]. Location-Centric [GDSB16]. Logic [SMM19]. Long [ACS19]. Long-Range [ACS19]. Low [Pet21, YSWZ18]. Low-Sampling-Rate [YSWZ18]. LSTM [STZ^{+}20].

Management [PBGA19, SS21]. Many [JS19]. Many-to-many [JS19]. Map [AFHW15, CFWW20, HBH^{+}21, JS19, Pet21, SKZ^{+}20, WK18, YSWZ18]. Map-Matching [CFWW20]. MapLUR [SKZ^{+}20]. Maps [BBS19, BVW16, KPS17, LNK^{+}21].

Markers [BVW16]. Markov [SMM19]. Massive [TDPT20]. Matches [CEGH17]. Matching [AWD^{+}18, BVW16, CFWW20, DPKW19, JS19, YSWZ18]. MaxCRS [AH17]. Maximum [CKM^{+}21]. MaxRS [AH17]. Measuring [AKAM17, CSKB19]. Medial [TIKG18]. Memory [FAMF16]. Mesh [TIKG18]. Meshes [FDMW20]. Method [MIF17]. Metrics [EP20]. Metropolitan [CSF^{+}19]. Microblogs [ZCLR16]. Minimal [Pet21]. Mining [LZLL20, PAB^{+}16, TKC17]. Mixture [ISNU17]. Mobile [BPM21, BDW18, CSF^{+}19, KKT^{+}18, LCKQ20]. Mobility [ARF19, CSF^{+}19, DHQ^{+}21, Gol19, QKZU20, SGBM20, SOdB^{+}20, DCAA21, WLL^{+}19, ZSFB20]. Mobility-on-Demand [ARF19]. Model [ARF19, DPG20, LA19, YÖR20]. Models [ISNU17]. Monitoring [AH17]. Most [PAB^{+}16]. Motion [YÖR20]. Movement [KTHK19]. Moving [MF15, TLF^{+}20]. Multi [TIKG18, WCC^{+}20, YÖR20]. Multi-Layered [TIKG18]. Multi-Pedestrian [YÖR20]. Multi-step [WCC^{+}20]. Multiflow [LA19]. Multiflow-Direction [LA19]. Multinomial [SMM19]. Multirow [GHN15]. Multivariate [LZLL20].

Natural [ABM16]. Navigation [DV21, TIKG18]. Near [HHRR20]. Near-accident [HHRR20]. Nearest [EEP16, HYL16, ZYW^{+}21]. Nearest-Neighbor [EEP16]. Neighbor [ABM16, EEP16, HYL16, ZYW^{+}21]. Nets [WCC^{+}20]. Network [DPG20, KPS17, WLL^{+}19]. Networks [BM20, BDW18, CSKB19, GDSB16, HHRR20, KTY^{+}18, MIF18, PK16, RHJC19, SMM19]. Number [LGLG19].

Obstructed [ZYW^{+}21]. occurrence [PAB^{+}16]. Occurrences [AKAM17].

Officer [STZ^{+}20]. Online [ZCLR16]. Open [BPS18]. Operational [QKZU20]. Optimal [KKT^{+}18, PWH21]. Optimizations [STZ^{+}20]. Optimizing [LCKQ20]. Our [PSTT16]. Outdoor [ABY17]. Overlay [BDKS19].

Paired [LZLL20]. Panorama [GHN15]. Papers [Are19, Are21]. Paradigm [SKZ^{+}20]. Parking [STZ^{+}20]. Participatory [MR21]. Particle [LH17]. Partition [LH17]. Partition-Based [LH17]. Partitioning [VBME21]. Passenger [MKW20]. Passenger-based [MKW20]. Passes [CEGH17]. Path [AFHW15, DPKW19]. Path-Based [AFHW15]. Paths [CFWW20, JS19]. Patrolling [STZ^{+}20]. Pattern [LZLL20]. Patterns [BASM21, PAB^{+}16, SOdB^{+}20]. Pebbles [BDKS19]. Pedestrian [TLF^{+}20, YÖR20]. People [ISNU17]. Personalized [AKM21, DV21, LGLG19, PK16, SSTN19]. Phenomena [BPM21]. Physically [CKM^{+}21]. Planning [FNCO20, MH19]. POI [SSTN19, ZWT^{+}20]. Point [BASM21, PBGA19, SOdB^{+}20]. POIs [MH19]. Pollution [SKZ^{+}20]. Polygon [CDFP21]. Polygonal [BMVS16]. Population [ISNU17, ZSFB20]. position [LNK^{+}21]. Post [DBB20]. Post-Disaster [DBB20]. Posterior [DPKW19]. Postprocessing [BVW16]. Practical [GHPS21]. Predictability [DCAA21]. Predicting [BPS18, BPM21]. Prediction [DPG20, LGLG19, WCC^{+}20]. Preferences [FNCO20]. Preserving [BMVS16]. Priority [MKW20]. Privacy [GKR16, NKTB20]. Privacy- [NKTB20]. Privacy-Aware [GKR16]. Probabilistic [Wer21]. Probe [LGLG19, PLHC19]. Problem [CDFP21, STZ^{+}20]. Problems [EP20, MICNC21]. Procedural [ZSBA21]. Process [BASM21]. Process-based [BASM21]. Processing [CCBS18, EEP16, HBH^{+}21, KS15, ZYW^{+}21]. Producing [BMNP16]. Products [EP20]. Programming [PWH21]. Projection [Pet21]. Protecting [GDSB16]. Providers [QKZU20]. Proximity [EP20, GDSB16, GHPS21]. Proximity-Based [GDSB16]. Public [HKK^{+}19].

Quality [LNK^{+}21]. Quality-of-position [LNK^{+}21]. Quantification [HKK^{+}19]. Queries [ATS^{+}16, CCBS18, DCY^{+}18, EEP16, GHPS21, MH19, NLC16, ZYW^{+}21]. Querying [KTY^{+}18].

Range [ACS19]. Rank [DCY^{+}18]. Raster [Wer21]. Rate [YSWZ18]. Raw [SSTN19]. Real [HHRR20]. Real-time [HHRR20]. Realistic [JSL^{+}20]. Recommendation [ZWT^{+}20]. Recommender [PK16]. Recurrent [DPG20]. ReFGeM [ZSFB20]. Region [WK18]. Regions [MF15]. Regression [SMM19]. regressive [ZGP19]. Regrets [CSKB19]. RegRocket [SMM19]. Regularity [DCAA21]. Relations [BDW18, SPKS16]. Release [NKTB20]. Relevance [DHQ^{+}21]. Relevant [EEP16]. Remote [DPG20]. Replication [HKK^{+}19]. Representations [PAB^{+}16]. Representative [ZSFB20]. Resolution [ACS19]. Resources [BPM21]. Restore [BMNP16]. Retrieval [MIF17]. Reverse [DCY^{+}18, HYL16]. Ride [CKT^{+}19, GKR16, JSL^{+}20]. Ride-Hailing [JSL^{+}20]. Ride-Sharing [CKT^{+}19]. Ridesharing [MH19]. Risk [LA19, RLA19]. Riso [SS21]. Riso-Tree [SS21]. Road [BM20, KPS17, KTY^{+}18, MIF18, PLHC19, RHJC19, WLL^{+}19]. Robust [BMNP16, DPKW19, KPS17]. Rocks [BDKS19]. Rounding [BMNP16]. Route [CN20, DPKW19, MIF17]. Routes [FHK^{+}18]. Routing [CSKB19, DV21, MICNC21, NT18].

Sampling [YSWZ18]. SARDINE [DPG20]. Scalable [SMM19, SS21]. Scale [LCKQ20, ZSBA21]. Scenes [ROOF17]. Scheduling [RHJC19]. Schematization [BMVS16, CDFP21]. SCPP [BASM21]. Search [AKM21, HYL16, TYZO15]. Section [Are21]. Selection [CN20, TYZO15]. Self [DPG20]. Self-Adaptive [DPG20]. Semi [ZWT^{+}20]. Semi-supervised [ZWT^{+}20]. Sensing [DPG20, HKK^{+}19, MR21]. Sensing-Assisted [HKK^{+}19]. Sensor [BPM21]. Sensors [BDKS19, TLF^{+}20]. Seq2Seq [WCC^{+}20]. SeqST [WCC^{+}20]. SeqST-GAN [WCC^{+}20]. Sequences [PWH21]. Series [LZLL20]. Server [TSK15]. Server-Assigned [TSK15]. Service [QKZU20]. Sharing [CKT^{+}19, GKR16]. Shortest [CFWW20, JS19]. Signal [HBH^{+}21, MKW20]. Signalized [MKW20]. Significance [AKAM17]. SIGSPATIAL [Are19, Are21]. Simplification [BMVS16]. Simulating [PSTT16]. Simulation [LH17]. Small [LGLG19]. Smart [DBB20]. SmarterROUTES [DV21]. Smoothing [BDKS19]. Snap [BMNP16]. Snapshots [MF15]. Social [AKM21, GDSB16, HYL16, PK16, YÖR20]. Solution [DV21]. Solutions [Pet21]. Solve [STZ^{+}20]. Solving [CDFP21]. sourced [HBH^{+}21]. Space [SOdB^{+}20, TKC17, ZYW^{+}21]. Space-Time [SOdB^{+}20]. Spaces [HKK^{+}19]. Spark [WMPH19]. Sparse [KTHK19]. Spatial [ATS^{+}16, AH17, BPS18, BASM21, CCBS18, CN20, DAGM21, FDMW20, LH17, MR21, PAB^{+}16, SPKS16, SS21, TSK15, TL20, VTSD18, VBME21, WMPH19, YKC20, ZSFB20, ZSBA21, ZCLR16, ZGP19]. Spatial-Textual [CCBS18]. Spatio [AWD^{+}18, BPM21, DPG20, WMPH19].

Spatio-Temporal

[AWD^{+}18, BPM21, DPG20, WMPH19]. Spatiotemporal [AKRH19, AKAM17, BDKS19, CEGH17, ISNU17, PAB^{+}16]. Special [Are19, Are21, Gol19, YKC20]. Spread [LH17]. Square [Pet21]. Stationarity [DCAA21]. step [WCC^{+}20]. Stepping [KTHK19]. Stochastic [ARF19]. Stone [KTHK19]. Stop [TKC17]. Stop-by [TKC17]. Storage [TPZ15]. Strategies [WLL^{+}19]. stream [HHRR20]. Streaming [AKM21, DAGM21]. Streams [AH17, BDKS19]. Street [AFHW15]. String [KTY^{+}18]. Structure [GHPS21, Wer21]. Subdivisions [BMVS16]. Subtrajectory [TDPT20]. supervised [ZWT^{+}20]. Supporting [GHPS21]. Surface [KS15]. Surrounds [BDW18]. SWARM [DAGM21]. Symbolic [GVD15]. Synthetic [WCM20]. Systems [ARF19, DAGM21, Gol19, PK16, SS21, YKC20].

Task [CN20, LCKQ20]. TDEFSI [WCM20]. Techniques [MR21]. Temporal [AKM21, AWD^{+}18, BPM21, DPG20, WMPH19]. Terrain [FAMF16, ROOF17]. Terrains [LA19, RLA19]. TerraNNI [ABM16]. Tetrahedral [FDMW20]. Textual [CCBS18]. Thematic [VTSD18]. Theory [WCM20]. Theory-guided [WCM20]. Throughput [LCKQ20]. Tilewise [KS15]. Time [EP20, LGLG19, LZLL20, SOdB^{+}20, HHRR20]. Top [CCBS18, PAB^{+}16]. Top- [CCBS18, PAB^{+}16]. Topological [PSD^{+}21, ZGP19]. Tracking [KPS17]. Tracks [FHK^{+}18]. Traffic [CSKB19, HHRR20]. Trajectories [CKM^{+}21, GHPS21, GVD15, KTHK19, KTY^{+}18, MAK^{+}18, MIF18, SSTN19, YSWZ18]. Trajectory [NKTB20, TPZ15, ZWT^{+}20]. Transit [BBS19, MKW20]. Transportation [AWD^{+}18]. Travel [LGLG19]. Treatment [EP20]. Tree [SS21]. Trees [FDMW20].

TRIFL [TPZ15]. Trip [MH19, ZWT^{+}20]. Trips [RHJC19]. Turbo [LCKQ20]. Turbo-GTS [LCKQ20]. Two [HHRR20]. Two-stream [HHRR20]. Type [BM20]. Type-based [BM20].

Understanding [CSF^{+}19, QKZU20, WLL^{+}19, ZWT^{+}20]. Unified [EP20]. Units [KS15]. Unordered [SMM19]. Urban [AWD^{+}18, FNCO20, Gol19, ROOF17, WCC^{+}20]. User [FNCO20]. Using [ABM16, BPS18, CFWW20, CEGH17, DPKW19, FHK^{+}18, HBH^{+}21, JSL^{+}20, LNK^{+}21, LGLG19, SMM19, SPKS16, SKZ^{+}20, TLF^{+}20, VBME21, ZSFB20, KTHK19].

Value [AKRH19]. Variables [SMM19]. Vector [TL20, WK18]. Vehicle [KPS17, MICNC21, PLHC19, YÖR20]. Vehicles [LGLG19]. Vehicular [FHK^{+}18]. Velocity [GDSB16]. Velocity-Based [GDSB16]. via [CSF^{+}19, ISNU17, KTY^{+}18, NLC16, TYZO15]. Video [HHRR20]. Viewshed [FAMF16]. Visibility [ROOF17]. Vision [TLF^{+}20]. Visited [SSTN19]. Visited-POI [SSTN19]. Visiting [BASM21]. Visual [LNK^{+}21]. Visualization [VTSD18]. Voronoi [ZYW^{+}21]. VRPDiv [MICNC21]. vs [PWH21].

Warping [EP20]. Watermarking [WK18]. Weighted [DCY^{+}18, DPKW19]. Wildfire [LH17].

References

Agarwal:2016:TNN

[ABM16] Pankaj K. Agarwal, Alex Beutel, and Thomas Mølhave. TerraNNI: Natural neighbor interpolation on 2D and 3D grids using a GPU. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2(2):7:1–7:31, July 2016. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL http://dl.acm.
org/citation.cfm?id=2786757.

Aly:2017:AEE

[ABY17] Heba Aly, Anas Basalamah, and Moustafa Youssef. Accurate and energy-efficient GPS-less outdoor localization. ACM Transactions on Spatial Algorithms and Systems (TSAS), 3(2):4:1–4:??, August 2017. CODEN ???? ISSN 2374-0353. URL https://

dl.acm.org/citation.cfm?id=
3085575.

Ayhan:2019:DDF

[ACS19] Samet Ayhan, Pablo Costas, and Hanan Samet. A data-driven framework for long-range aircraft conflict detection and resolution. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(4):24:1–24:??, December 2019. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3328832.

Ahmed:2015:PBD

[AFHW15] Mahmuda Ahmed, Brittany Terese Fasy, Kyle S. Hickmann, and Carola Wenk. A path-based distance for street map comparison. ACM Transactions on Spatial Algorithms and Systems (TSAS), 1(1):3:1–3:28, August 2015. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL http://dl.acm.
org/citation.cfm?id=2729977.

Amagata:2017:GFM

[AH17] Daichi Amagata and Takahiro Hara. A general framework for MaxRS and MaxCRS monitoring in spatial data streams. ACM Transactions on Spatial Algorithms and Systems (TSAS), 3 (1):1:1–1:34, May 2017. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic). URL

http://dl.acm.org/citation.
cfm?id=3080554.

Aydin:2017:MSS

[AKAM17] Berkay Aydin, Ahmet Kucuk, Rafal A. Angryk, and Petrus C. Martens. Measuring the significance of spatiotemporal cooccurrences. ACM Transactions on Spatial Algorithms and Systems (TSAS), 3(3):9:1–9:??, November 2017. CODEN ???? ISSN 2374-0353. URL https://

dl.acm.org/citation.cfm?id=
3139351.

Almaslukh:2021:TGS

[AKM21] Abdulaziz Almaslukh, Yunfan Kang, and Amr Magdy. Temporal geo-social personalized keyword search over streaming data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7(4):20:1–20:28, December 2021. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3473006.

Aly:2019:BBC

[AKRH19] Heba Aly, John Krumm, Gireeja Ranade, and Eric Horvitz. To buy or not to buy: Computing value of spatiotemporal information. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(4):22:1–22:??, December 2019. CODEN ???? ISSN 2374-0353. URL https://

dl.acm.org/citation.cfm?id=
3320431.

Aref:2019:ISI

[Are19] Walid G. Aref. Introduction to the special issue on the best papers from the 2017 ACM SIGSPATIAL Conference. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5 (1):1:1–1:??, June 2019. CO-DEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3325134.

Aref:2021:ISS

[Are21] Walid G. Aref. Introduction to the special section on the best papers from the 2019 ACM SIGSPATIAL Conference. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7 (4):16:1–16:2, December 2021. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3485049.

Albert:2019:IMD

[ARF19] Marc Albert, Claudio Ruch, and Emilio Frazzoli. Imbalance in mobility-on-demand systems: A stochastic model and distributed control approach. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(2):13:1– 13:??, August 2019. CODEN ????

ISSN 2374-0353. URL https://

dl.acm.org/citation.cfm?id=
3325914.

Ali:2016:SCQ

[ATS^{+}16] Mohammed Eunus Ali, Egemen Tanin, Peter Scheuermann, Sarana Nutanong, and Lars Kulik. Spatial consensus queries in a collaborative environment. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2 (1):3:1–3:37, April 2016. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic). URL

http://dl.acm.org/citation.
cfm?id=2829943.

Ayala:2018:STM

[AWD^{+}18] Daniel Ayala, Ouri Wolfson, Bhaskar Dasgupta, Jie Lin, and Bo Xu. Spatio-temporal matching for urban transportation applications. ACM Transactions on Spatial Algorithms and Systems (TSAS), 3(4):11:1–11:??, May 2018. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3183344.

Brito:2021:SPP

[BASM21] Denise E. F. Brito, Renato M.

Assunção, Roberto C. S. N. P. Souza, and Wagner Meira, Jr. SCPP: a point process-based clustering of spatial visiting patterns. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7(1):5:1–5:30, January 2021. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3423405.

Bast:2019:EGG

[BBS19] Hannah Bast, Patrick Brosi, and Sabine Storandt. Efficient generation of geographically accurate transit maps. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5 (4):25:1–25:??, December 2019. CODEN ???? ISSN 2374-

                                        0353. URL https://dl.acm.
                                        org/citation.cfm?id=3337790.

Brown:2019:RPS

[BDKS19] Philip E. Brown, Tamraparni Dasu, Yaron Kanza, and Divesh Srivastava. From rocks to pebbles: Smoothing spatiotemporal data streams in an overlay of sensors. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(3):18:1–18:??, September 2019. CODEN ???? ISSN 2374-0353. URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3329677.

Both:2018:ISE

[BDW18] Alan Both, Matt Duckham, and Michael F. Worboys. Identifying surrounds and engulfs relations in mobile and coordinate-free geosensor networks. ACM Transactions on Spatial Algorithms and Systems (TSAS), 4(2):6:1– 6:??, August 2018. CODEN ???? ISSN 2374-0353. URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3234505.

Babu:2020:TTB

[BM20] Sarath Babu and B. S. Manoj. Toward a type-based analysis of road networks. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6(4):28:1– 28:45, August 2020. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic). URL

https://dl.acm.org/doi/10.
1145/3397579.

Belussi:2016:SRR

[BMNP16] Alberto Belussi, Sara Migliorini, Mauro Negri, and Giuseppe Pelagatti. Snap rounding with restore: An algorithm for producing robust geometric datasets. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2 (1):1:1–1:36, April 2016. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic). URL

http://dl.acm.org/citation.
cfm?id=2811256.

Buchin:2016:APS

[BMVS16] Kevin Buchin, Wouter Meulemans, André Van Renssen, and Bettina Speckmann. Areapreserving simplification and schematization of polygonal subdivisions. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2(1):2:1–2:36, April 2016. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL http://dl.acm.
org/citation.cfm?id=2818373.

Bose:2021:PST

[BPM21] Sunanda Bose, Sumit Kumar Paul, and Nandini Mukherjee. Predicting spatio-temporal phenomena of mobile resources in sensor cloud infrastructure. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7 (3):11:1–11:38, September 2021. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3446936.

Belesiotis:2018:APS

[BPS18] Alexandros Belesiotis, George Papadakis, and Dimitrios Skoutas. Analyzing and predicting spatial crime distribution using crowdsourced and open data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 3(4):12:1–12:??, May 2018. CODEN ???? ISSN 2374-

                                        0353. URL https://dl.acm.
                                        org/citation.cfm?id=3190345.

Budig:2016:MLM

[BVW16] Benedikt Budig, Thomas C. Van Dijk, and Alexander Wolff. Matching labels and markers in historical maps: An algorithm with interactive postprocessing. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2 (4):13:1–13:24, November 2016. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

                                        tronic). URL http://dl.acm.
                                        org/citation.cfm?id=2994598.

Choudhury:2018:BPT

[CCBS18] Farhana M. Choudhury, J. Shane Culpepper, Zhifeng Bao, and Timos Sellis. Batch processing of top-k spatial-textual queries. ACM Transactions on Spatial Algorithms and Systems (TSAS), 3 (4):13:1–13:??, May 2018. CO-DEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3196155.

Cicerone:2021:CPS

[CDFP21] Serafino Cicerone, Mattia D’emidio, Daniele Frigioni, and Filippo Tirabassi Pascucci. Combining polygon schematization and decomposition approaches for solving the cavity decomposition problem. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7 (4):22:1–22:37, June 2021. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3462760.

Chawla:2017:CPF

[CEGH17] Sanjay Chawla, Joël Estephan, Joachim Gudmundsson, and Michael Horton. Classification of passes in football matches using spatiotemporal data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 3 (2):6:1–6:??, August 2017. CO-DEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3105576.

Chambers:2020:MMU

[CFWW20] Erin Chambers, Brittany Terese Fasy, Yusu Wang, and Carola Wenk. Map-matching using shortest paths. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6(1):6:1– 6:17, February 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        abs/10.1145/3368617.

Custers:2021:MPC

[CKM^{+}21] Bram Custers, Mees Van De Kerkhof, Wouter Meulemans, Bettina Speckmann, and Frank Staals. Maximum physically consistent trajectories. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7 (4):17:1–17:33, June 2021. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3452378.

Correa:2019:CAR

[CKT^{+}19] Oscar Correa, A. K. M. Mustafizur Rahman Khan, Egemen Tanin, Lars Kulik, and Kotagiri Ramamohanarao. Congestion-aware ridesharing. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(1):5:1–5:??, June 2019. CODEN ???? ISSN 2374-

                                        0353. URL https://dl.acm.
                                        org/citation.cfm?id=3317639.

Costa:2020:RTS

[CN20] Camila F. Costa and Mario A. Nascimento. In-route task selection in spatial crowdsourcing. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (2):7:1–7:45, February 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3368268.

Cao:2019:UMC

[CSF^{+}19] Hancheng Cao, Jagan Sankaranarayanan, Jie Feng, Yong Li, and Hanan Samet. Understanding metropolitan crowd mobility via mobile cellular accessing data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(2):8:1–8:??, August 2019. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3323345.

Cabannes:2019:RRN

[CSKB19] Théophile Cabannes, Marco Sangiovanni, Alexander Keimer, and Alexandre M. Bayen. Regrets in routing networks: Measuring the impact of routing apps in traffic. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(2):9:1–9:??, August 2019. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3325916.

Daghistani:2021:SAL

[DAGM21] Anas Daghistani, Walid G. Aref, Arif Ghafoor, and Ahmed R. Mahmood. SWARM: Adaptive load balancing in distributed streaming systems for big spatial data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7(3):14:1– 14:43, September 2021. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3460013.

Das:2020:EDD

[DBB20] Nabanita Das, Souvik Basu, and Sipra Das Bit. Efficient Drop-Box deployment toward improving post-disaster information exchange in a smart city. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (2):9:1–9:18, February 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3373645.

Teixeira:2021:ISR

[DCAA21] Douglas Do Couto Teixeira, Aline Carneiro Viana, Jussara M. Almeida, and Mário S. Alvim. The impact of stationarity, regularity, and context on the predictability of individual human mobility. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7(4):19:1–19:24, June 2021. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3459625.

Dong:2018:WAR

[DCY^{+}18] Yuyang Dong, Hanxiong Chen, Jeffrey Xu Yu, Kazutaka Furuse, and Hiroyuki Kitagawa. Weighted aggregate reverse rank queries. ACM Transactions on Spatial Algorithms and Systems (TSAS), 4(2):5:1–5:??, August 2018. CODEN ???? ISSN 2374-

                                        0353. URL https://dl.acm.
                                        org/citation.cfm?id=3225216.

Damiani:2021:LRD

[DHQ^{+}21] Maria Luisa Damiani, Fatima Hachem, Christian Quadri, Matteo Rossini, and Sabrina Gaito. On location relevance and diversity in human mobility data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7 (2):7:1–7:38, February 2021. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3423404.

Das:2020:SSA

[DPG20] Monidipa Das, Mahardhika Pratama, and Soumya K. Ghosh. SARDINE: a self-adaptive recurrent deep incremental network model for spatio-temporal prediction of remote sensing data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (3):16:1–16:26, May 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3380972.

Doocy:2019:RPM

[DPKW19] Lauren Doocy, Steven D. Prager, Joseph T. Kider, Jr., and R. Paul Wiegand. Robust path matching and anomalous route detection using posterior weighted graphs. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(2):14:1–14:??, August 2019. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3338905.

DeBock:2021:SDD

[DV21] Jelle De Bock and Steven Verstockt. SmarterROUTES — a data-driven context-aware solution for personalized dynamic routing and navigation. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7 (1):2:1–2:25, January 2021. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3402125.

Efstathiades:2016:EPR

[EEP16] Christodoulos Efstathiades, Alexandros Efentakis, and Dieter Pfoser. Efficient processing of relevant nearest-neighbor queries. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2 (3):9:1–9:28, October 2016. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

                                        tronic). URL http://dl.acm.
                                        org/citation.cfm?id=2934675.

Emiris:2020:PEM

[EP20] Ioannis Z. Emiris and Ioannis Psarros. Products of Euclidean metrics, applied to proximity problems among curves: Unified treatment of discrete Fréchet and dynamic time warping distances. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (4):27:1–27:20, August 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3397518.

Ferreira:2016:EEM

[FAMF16] Chaulio R. Ferreira, Marcus V. A. Andrade, Salles V. G. Magalhães, and W. Randolph Franklin. An efficient external memory algorithm for terrain viewshed computation. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2 (2):6:1–6:17, July 2016. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic). URL

http://dl.acm.org/citation.
cfm?id=2903206.

Fellegara:2020:TTF

[FDMW20] Riccardo Fellegara, Leila De Floriani, Paola Magillo, and Kenneth Weiss. Tetrahedral trees: a family of hierarchical spatial indexes for tetrahedral meshes. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (4):23:1–23:34, August 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3385851.

Fujino:2018:DDI

[FHK^{+}18] Takumi Fujino, Atsushi Hashimoto, Hidekazu Kasahara, Mikihiko Mori, Masaaki Iiyama, and Michihiko Minoh. Detecting deviations from intended routes using vehicular GPS tracks. ACM Transactions on Spatial Algorithms and Systems (TSAS), 4 (1):1:1–1:??, June 2018. CO-DEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3204455.

Ferreira:2020:DLA

[FNCO20] Danielle L. Ferreira, Bruno A. A.

Nunes, Carlos Alberto V. Campos, and Katia Obraczka. A deep learning approach for identifying user communities based on geographical preferences and its applications to urban and environmental planning. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (3):17:1–17:24, May 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3380970.

Ghinita:2016:PAV

[GDSB16] Gabriel Ghinita, Maria Luisa Damiani, Claudio Silvestri, and Elisa Bertino. Protecting against velocity-based, proximity-based, and external event attacks in location-centric social networks. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2 (2):8:1–8:36, July 2016. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic). URL

http://dl.acm.org/citation.
cfm?id=2910580.

Gemsa:2015:MBL

[GHN15] Andreas Gemsa, Jan-Henrik Haunert, and Martin Nöllenburg. Multirow boundary-labeling algorithms for panorama images. ACM Transactions on Spatial Algorithms and Systems (TSAS), 1 (1):1:1–1:30, August 2015. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL http://dl.acm.
org/citation.cfm?id=2794299.

Gudmundsson:2021:PIS

[GHPS21] Joachim Gudmundsson, Michael Horton, John Pfeifer, and Martin P. Seybold. A practical index structure supporting Fréchet proximity queries among trajectories. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7(3):15:1– 15:33, September 2021. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3460121.

Goel:2016:PAD

[GKR16] Preeti Goel, Lars Kulik, and Kotagiri Ramamohanarao. Privacyaware dynamic ride sharing. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2 (1):4:1–4:41, April 2016. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic). URL

http://dl.acm.org/citation.
cfm?id=2845080.

Gollapudi:2019:ISI

[Gol19] Sreenivas Gollapudi. Introduction to the special issue on urban mobility: Algorithms and systems. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(2):7:1–7:??, August 2019. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3346023.

Guting:2015:ST

[GVD15] Ralf Hartmut Güting, Fabio Valdés, and Maria Luisa Damiani. Symbolic trajectories. ACM Transactions on Spatial Algorithms and Systems (TSAS), 1(2):7:1–7:51, November 2015. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL http://dl.acm.
org/citation.cfm?id=2786756.

He:2021:MIA

[HBH^{+}21] Eric He, Fan Bai, Curtis Hay, Jinzhu Chen, and Vijayakumar Bhagavatula. A map inference approach using signal processing from crowd-sourced GPS data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7 (2):9:1–9:23, February 2021. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3431785.

Huang:2020:IIT

[HHRR20] Xiaohui Huang, Pan He, Anand Rangarajan, and Sanjay Ranka. Intelligent intersection: Twostream convolutional networks for real-time near-accident detection in traffic video. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (2):10:1–10:28, February 2020. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3373647.

Hemminki:2019:CRS

[HKK^{+}19] Samuli Hemminki, Keisuke Kuribayashi, Shin’ichi Konomi, Petteri Nurmi, and Sasu Tarkoma. Crowd replication: Sensingassisted quantification of human behavior in public spaces. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5 (3):15:1–15:??, September 2019. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3317666.

Hung:2016:SIA

[HYL16] Hui-Ju Hung, De-Nian Yang, and Wang-Chien Lee. Social influence-aware reverse nearest neighbor search. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2(3):12:1– 12:35, October 2016. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic). URL

http://dl.acm.org/citation.
cfm?id=2964906.

Iwata:2017:EPF

[ISNU17] Tomoharu Iwata, Hitoshi Shimizu, Futoshi Naya, and Naonori Ueda. Estimating people flow from spatiotemporal population data via collective graphical mixture models. ACM Transactions on Spatial Algorithms and Systems (TSAS), 3(1):2:1–2:18, May 2017. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL http://dl.acm.
org/citation.cfm?id=3080555.

Jagadeesh:2019:FCC

[JS19] George R. Jagadeesh and Thambipillai Srikanthan. Fast computation of clustered many-to-many shortest paths and its application to map matching. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5 (3):17:1–17:??, September 2019. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3329676.

Jauhri:2020:GRR

[JSL^{+}20] Abhinav Jauhri, Brad Stocks, Jian Hui Li, Koichi Yamada, and John Paul Shen. Generating realistic ride-hailing datasets using GANs. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6(3):18:1–18:14, May 2020. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3380968.

Khan:2018:ECO

[KKT^{+}18] A. K. M. Mustafizur Rahman Khan, Lars Kulik, Egemen Tanin, Hua Hua, and Tanzima Hashem. Efficient computation of the optimal accessible location for a group of mobile agents. ACM Transactions on Spatial Algorithms and Systems (TSAS), 4 (4):10:1–10:??, October 2018. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3239124.

Karagiorgou:2017:LAM

[KPS17] Sophia Karagiorgou, Dieter Pfoser, and Dimitrios Skoutas. A layered approach for more robust generation of road network maps from vehicle tracking data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 3 (1):3:1–3:21, May 2017. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic). URL

http://dl.acm.org/citation.
cfm?id=3061713.

Kovanen:2015:TAC

[KS15] Janne Kovanen and Tapani Sarjakoski. Tilewise accumulated cost surface computation with graphics processing units. ACM Transactions on Spatial Algorithms and Systems (TSAS), 1(2):8:1–8:27, November 2015. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL http://dl.acm.
org/citation.cfm?id=2803172.

Kannangara:2019:SSG

[KTHK19] Sameera Kannangara, Egemen Tanin, Aaron Harwood, and Shanika Karunasekera. Stepping stone graph: A graph for finding movement corridors using sparse trajectories. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5 (4):23:1–23:??, December 2019. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3324883.

Koide:2018:EIQ

[KTY^{+}18] Satoshi Koide, Yukihiro Tadokoro, Takayoshi Yoshimura, Chuan Xiao, and Yoshiharu Ishikawa. Enhanced indexing and querying of trajectories in road networks via string algorithms. ACM Transactions on Spatial Algorithms and Systems (TSAS), 4(1):3:1–3:??, June 2018. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3200200.

Lowe:2019:FRA

[LA19] Aaron Lowe and Pankaj K. Agarwal. Flood-risk analysis on terrains under the multiflowdirection model. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(4):26:1–26:??, December 2019. CODEN ???? ISSN 2374-0353. URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3340707.

Li:2020:TGF

[LCKQ20] Wei Li, Haiquan Chen, Wei-Shinn Ku, and Xiao Qin. Turbo-GTS: a fast framework of optimizing task throughput for largescale mobile crowdsourcing. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (1):1:1–1:29, February 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        abs/10.1145/3363450.

Li:2019:PTT

[LGLG19] Yang Li, Dimitrios Gunopulos, Cewu Lu, and Leonidas J. Guibas. Personalized travel time prediction using a small number of probe vehicles. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(1):4:1– 4:??, June 2019. CODEN ???? ISSN 2374-0353. URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3317663.

Long:2017:SPB

[LH17] Yuan Long and Xiaolin Hu. Spatial partition-based particle filtering for data assimilation in wildfire spread simulation. ACM Transactions on Spatial Algorithms and Systems (TSAS), 3(2):5:1–5:??, August 2017. CODEN ???? ISSN 2374-

                                        0353. URL https://dl.acm.
                                        org/citation.cfm?id=3099471.

Laoudias:2021:IQP

[LNK^{+}21] Christos Laoudias, Artyom Nikitin, Panagiotis Karras, Moustafa Youssef, and Demetrios Zeinalipour-Yazti. Indoor quality-of-position visual assessment using crowdsourced fingerprint maps. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7 (2):10:1–10:32, February 2021. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3433026.

Li:2020:CPM

[LZLL20] Qingzhe Li, Liang Zhao, Yi-Ching Lee, and Jessica Lin. Contrast pattern mining in paired multivariate time series of a controlled driving behavior experiment. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6(4):25:1–25:28, August 2020. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3397272.

Mahmood:2018:DBI

[MAK^{+}18] Ahmed R. Mahmood, Ahmed M.

Aly, Tatiana Kuznetsova, Saleh Basalamah, and Walid G. Aref. Disk-based indexing of recent trajectories. ACM Transactions on Spatial Algorithms and Systems (TSAS), 4(3):7:1–7:??, September 2018. CODEN ???? ISSN 2374-0353. URL https://

dl.acm.org/citation.cfm?id=
3234941.

Mckennney:2015:GMR

[MF15] Mark Mckennney and Roger Frye. Generating moving regions from snapshots of complex regions. ACM Transactions on Spatial Algorithms and Systems (TSAS), 1(1):4:1–4:30, August 2015. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

                                        tronic). URL http://dl.acm.
                                        org/citation.cfm?id=2774220.

Mahin:2019:AAR

[MH19] Mehnaz Tabassum Mahin and Tanzima Hashem. Activity-aware ridesharing group trip planning queries for flexible POIs. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5 (3):20:1–20:??, September 2019. CODEN ???? ISSN 2374-

                                        0353. URL https://dl.acm.
                                        org/citation.cfm?id=3341818.

Mariescu-Istodor:2021:VDC

[MICNC21] Radu Mariescu-Istodor, Alexandru Cristian, Mihai Negrea, and Peiwei Cao. VRPDiv: a divide and conquer framework for large vehicle routing problems. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7(4):23:1–23:41, December 2021. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3474832.

Mariescu-Istodor:2017:GBM

[MIF17] Radu Mariescu-Istodor and Pasi Fränti. Grid-based method for GPS route analysis for retrieval.

ACM Transactions on Spatial Algorithms and Systems (TSAS), 3(3):8:1–8:??, November 2017. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3125634.

Mariescu-Istodor:2018:CIR

[MIF18] Radu Mariescu-Istodor and Pasi Fränti. CellNet: Inferring road networks from GPS trajectories. ACM Transactions on Spatial Algorithms and Systems (TSAS), 4(3):8:1–8:??, September 2018. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3234692.

Mishra:2020:TSP

[MKW20] Suman Mishra, Lina Kattan, and S. C. Wirasinghe. Transit signal priority along a signalized arterial: a passenger-based approach. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (1):5:1–5:19, February 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
abs/10.1145/3355611.

Middya:2021:SIT

[MR21] Asif Iqbal Middya and Sarbani Roy. Spatial interpolation techniques on participatory sensing data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7(3):13:1– 13:32, September 2021. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3457609.

Naghizade:2020:PCA

[NKTB20] Elham Naghizade, Lars Kulik, Egemen Tanin, and James Bailey. Privacy- and context-aware release of trajectory data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (1):3:1–3:25, February 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        abs/10.1145/3363449.

Niu:2016:LED

[NLC16] Wei Niu, Zhijiao Liu, and James Caverlee. On local expert discovery via geo-located crowds, queries, and candidates. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2 (4):14:1–14:24, November 2016. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

                                        tronic). URL http://dl.acm.
                                        org/citation.cfm?id=2994599.

Nur:2018:GRI

[NT18] Abdullah Yasin Nur and Mehmet Engin Tozal. Geography and routing in the Internet. ACM Transactions on Spatial Algorithms and Systems (TSAS), 4(4):11:1–11:??, October 2018. CODEN ???? ISSN 2374-0353. URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3239162.

Pillai:2016:MMT

[PAB^{+}16] Karthik Ganesan Pillai, Rafal A. Angryk, Juan M. Banda, Dustin Kempton, Berkay Aydin, and Petrus C. Martens. Mining at most top-K % spatiotemporal co-occurrence patterns in datasets with extended spatial representations. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2(3):10:1– 10:27, October 2016. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic). URL

http://dl.acm.org/citation.
cfm?id=2936775.

Pavlovic:2019:DCP

[PBGA19] Mirjana Pavlovic, Kai-Niklas Bastian, Hinnerk Gildhoff, and Anastasia Ailamaki. Dictionary compression in point cloud data management. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(1):3:1–3:??, June 2019. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3299770.

Petroff:2021:SEA

[Pet21] Matthew A. Petroff. A square equal-area map projection with low angular distortion, minimal cusps, and closed-form solutions. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7(4):21:1–21:16, December 2021. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3460521.

Purushotham:2016:PGR

[PK16] Sanjay Purushotham and C.- C. Jay Kuo. Personalized group recommender systems for location- and event-based social networks. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2(4):16:1–16:29, November 2016. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic). URL http://

                                        dl.acm.org/citation.cfm?id=
                                        2987381.

Pietrobon:2019:ARC

[PLHC19] Davide Pietrobon, Andrew P. Lewis, and Gavin S. Heverly-Coulson. An algorithm for road closure detection from vehicle probe data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(2):12:1–12:??, August 2019. CODEN ???? ISSN 2374-0353. URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3325912.

Paoluzzi:2021:TCA

[PSD^{+}21] Alberto Paoluzzi, Vadim Shapiro, Antonio Dicarlo, Francesco Furiani, Giulio Martella, and Giorgio Scorzelli. Topological computing of arrangements with (co)chains. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7 (1):1:1–1:29, January 2021. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3401988.

Pelekis:2016:SOL

[PSTT16] Nikos Pelekis, Stylianos Sideridis, Panagiotis Tampakis, and Yannis Theodoridis. Simulating our LifeSteps by example. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2 (3):11:1–11:39, October 2016.

CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL http://dl.acm.
org/citation.cfm?id=2937753.

Peng:2021:FOS

[PWH21] Dongliang Peng, Alexander Wolff, and Jan-Henrik Haunert. Finding optimal sequences for area aggregation — A^{⋆} vs. integer linear programming. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7 (1):4:1–4:40, January 2021. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3409290.

Qian:2020:UOD

[QKZU20] Xinwu Qian, Dheeraj Kumar, Wenbo Zhang, and Satish V. Ukkusuri. Understanding the operational dynamics of mobility service providers: a case of Uber. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6(2):12:1–12:20, February 2020. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3378888.

Rayhan:2019:ESG

[RHJC19] Yeasir Rayhan, Tanzima Hashem, Roksana Jahan, and Muhammad Aamir Cheema. Efficient scheduling of generalized group trips in road networks. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(2):10:1–10:??, August 2019. CODEN ???? ISSN 2374-

                                        0353. URL https://dl.acm.
                                        org/citation.cfm?id=3325915.

Rav:2019:FRA

[RLA19] Mathias Rav, Aaron Lowe, and Pankaj K. Agarwal. Flood risk analysis on terrains. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(1):2:1– 2:??, June 2019. CODEN ???? ISSN 2374-0353. URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3295459.

Robles-Ortega:2017:EVD

[ROOF17] M. D. Robles-Ortega, L. Ortega, and F. R. Feito. Efficient visibility determination in urban scenes considering terrain information. ACM Transactions on Spatial Algorithms and Systems (TSAS), 3(3):10:1–10:??, November 2017. CODEN ???? ISSN 2374-

                                        0353. URL https://dl.acm.
                                        org/citation.cfm?id=3152536.

Smolyak:2020:CIG

[SGBM20] Daniel Smolyak, Kathryn Gray, Sarkhan Badirli, and George Mohler. Coupled IGMM-GANs with applications to anomaly detection in human mobility data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (4):24:1–24:14, August 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3385809.

Steininger:2020:MEN

[SKZ^{+}20] Michael Steininger, Konstantin Kobs, Albin Zehe, Florian Lautenschlager, Martin Becker, and Andreas Hotho. MapLUR: Exploring a new paradigm for estimating air pollution using deep learning on map images. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (3):19:1–19:24, May 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3380973.

Sabek:2019:RSM

[SMM19] Ibrahim Sabek, Mashaal Musleh, and Mohamed F. Mokbel. RegRocket: Scalable multinomial autologistic regression with unordered categorical variables using Markov logic networks. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5 (4):27:1–27:??, December 2019. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3366459.

Souza:2020:STD

[SOdB^{+}20] Roberto C. S. N. P. Souza, Derick M. Oliveira, Denise E. F. de Brito, Renato M. Assunção, and Wagner Meira Jr. Space-time drift point detection in mobility patterns. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6(1):4:1–4:24, February 2020. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL https://dl.acm.
org/doi/abs/10.1145/3360721.

Skoumas:2016:LEU

[SPKS16] Georgios Skoumas, Dieter Pfoser, Anastasios Kyrillidis, and Timos Sellis. Location estimation using crowdsourced spatial relations. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2(2):5:1–5:23, July 2016. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

                                        tronic). URL http://dl.acm.
                                        org/citation.cfm?id=2894745.

Sun:2021:RTE

[SS21] Yuhan Sun and Mohamed Sarwat. Riso-Tree: an efficient and scalable index for spatial entities in graph database management systems. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7(3):12:1– 12:39, September 2021. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3450945.

Suzuki:2019:PVP

[SSTN19] Jun Suzuki, Yoshihiko Suhara, Hiroyuki Toda, and Kyosuke Nishida. Personalized visited-POI assignment to individual raw GPS trajectories. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5 (3):16:1–16:??, September 2019. CODEN ???? ISSN 2374-

                                        0353. URL https://dl.acm.
                                        org/citation.cfm?id=3317667.

Shao:2020:ILA

[STZ^{+}20] Wei Shao, Siyu Tan, Sichen Zhao, Kyle Kai Qin, Xinhong Hei, Jeffrey Chan, and Flora D. Salim. Incorporating LSTM auto-encoders in optimizations to solve parking officer patrolling problem. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6(3):20:1–20:21, May 2020. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3380966.

Tampakis:2020:DSJ

[TDPT20] Panagiotis Tampakis, Christos Doulkeridis, Nikos Pelekis, and Yannis Theodoridis. Distributed subtrajectory join on massive datasets. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6(2):8:1–8:29, February 2020. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3373642.

Toll:2018:MAM

[TIKG18] Wouter Van Toll, Atlas F. Cook Iv, Marc J. Van Kreveld, and Roland Geraerts. The medial axis of a multi-layered environment and its application as a navigation mesh. ACM Transactions on Spatial Algorithms and Systems (TSAS), 4(1):2:1–2:??, June 2018. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3204456.

Teng:2017:TMS

[TKC17] Shan-Yun Teng, Wei-Shinn Ku, and Kun-Ta Chuang. Toward mining stop-by behaviors in indoor space. ACM Transactions on Spatial Algorithms and Systems (TSAS), 3(2):7:1–7:??, August 2017. CODEN ????

ISSN 2374-0353. URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3106736.

Touya:2020:DLE

[TL20] Guillaume Touya and Imran Lokhat. Deep learning for enrichment of vector spatial databases: Application to highway interchange. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6(3):21:1–21:21, May 2020. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3382080.

Tokuda:2020:NAP

[TLF^{+}20] Eric K. Tokuda, Yitzchak Lockerman, Gabriel B. A. Ferreira, Ethan Sorrelgreen, David Boyle, Roberto M. Cesar, Jr., and Claudio T. Silva. A new approach for pedestrian density estimation using moving sensors and computer vision. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6(4):26:1–26:20, August 2020. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3397575.

That:2015:TGT

[TPZ15] Dai Hai Ton That, Iulian Sandu Popa, and Karine Zeitouni. TRIFL: A generic trajectory index for flash storage. ACM Transactions on Spatial Algorithms and Systems (TSAS), 1(2):6:1–6:44, November 2015. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL http://dl.acm.
org/citation.cfm?id=2786758.

To:2015:SAS

[TSK15] Hien To, Cyrus Shahabi, and Leyla Kazemi. A server-assigned spatial crowdsourcing framework. ACM Transactions on Spatial Algorithms and Systems (TSAS), 1(1):2:1–2:28, August 2015. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL http://dl.acm.
org/citation.cfm?id=2729713.

Tang:2015:EGF

[TYZO15] Suhua Tang, Yi Yu, Roger Zimmermann, and Sadao Obana. Efficient geo-fencing via hybrid hashing: A combination of bucket selection and in-bucket binary search. ACM Transactions on Spatial Algorithms and Systems (TSAS), 1(2):5:1–5:22, November 2015. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL http://dl.acm.
org/citation.cfm?id=2774219.

Vu:2021:UDL

[VBME21] Tin Vu, Alberto Belussi, Sara Migliorini, and Ahmed Eldway. Using deep learning for big spatial data partitioning. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7 (1):3:1–3:37, January 2021. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3402126.

Vollmer:2018:HSA

[VTSD18] Jan Ole Vollmer, Matthias Trapp, Heidrun Schumann, and Jürgen Döllner. Hierarchical spatial aggregation for level-of-detail visualization of 3D thematic data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 4(3):9:1–9:??, September 2018. CODEN ???? ISSN 2374-0353. URL https://

dl.acm.org/citation.cfm?id=
3234506.

Wang:2020:SGS

[WCC^{+}20] Senzhang Wang, Jiannong Cao, Hao Chen, Hao Peng, and Zhiqiu Huang. SeqST-GAN: Seq2Seq generative adversarial nets for multi-step urban crowd flow prediction. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6(4):22:1–22:24, August 2020. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3378889.

Wang:2020:TTG

[WCM20] Lijing Wang, Jiangzhuo Chen, and Madhav Marathe. TDEFSI: Theory-guided deep learningbased epidemic forecasting with synthetic information. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (3):15:1–15:39, May 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3380971.

Werner:2021:GAE

[Wer21]  Martin Werner.  GloBiMap-

sAI: an AI-enhanced probabilistic data structure for global raster datasets. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7(4):18:1–18:24, June 2021. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3453184.

Wang:2018:VMF

[WK18] Nana Wang and Mohan Kankanhalli. 2D vector map fragile watermarking with region location. ACM Transactions on Spatial Algorithms and Systems (TSAS), 4(4):12:1–12:??, October 2018. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3239163.

Wang:2019:ADR

[WLL^{+}19] Haiquan Wang, Yilin Li, Guoping Liu, Xiang Wen, and Xiaohu Qie. Accurate detection of road network anomaly by understanding crowd’s driving strategies from human mobility. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(2):11:1–11:??, August 2019. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3325913.

Whitman:2019:DSS

[WMPH19] Randall T. Whitman, Bryan G.

Marsh, Michael B. Park, and Erik G. Hoel. Distributed spatial and spatio-temporal join on Apache spark. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5(1):6:1– 6:??, June 2019. CODEN ???? ISSN 2374-0353. URL https://

dl.acm.org/citation.cfm?id=
3325135.

Youssef:2020:ISI

[YKC20] Moustafa Youssef, John Krum, and Muhammad Aamir Cheema. Introduction to the special issue on deep learning for spatial algorithms and systems. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (3):14:1–14:2, May 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3386878.

Yang:2020:SFB

[YÖR20] Dongfang Yang, Ümit Özgüner, and Keith Redmill. A social force based pedestrian motion model considering multi-pedestrian interaction with a vehicle. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6 (2):11:1–11:27, February 2020. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3373646.

Yin:2018:FBM

[YSWZ18] Yifang Yin, Rajiv Ratn Shah, Guanfeng Wang, and Roger Zimmermann. Feature-based map matching for low-sampling-rate GPS trajectories. ACM Transactions on Spatial Algorithms and Systems (TSAS), 4(2):4:1– 4:??, August 2018. CODEN ???? ISSN 2374-0353. URL https://

dl.acm.org/citation.cfm?id=
3223049.

Zhao:2016:OSE

[ZCLR16] Liang Zhao, Feng Chen, Chang-Tien Lu, and Naren Ramakrishnan. Online spatial event forecasting in microblogs. ACM Transactions on Spatial Algorithms and Systems (TSAS), 2 (4):15:1–15:39, November 2016. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

tronic). URL http://dl.acm.
org/citation.cfm?id=2997642.

Zhao:2019:SAR

[ZGP19] Liang Zhao, Olga Gkountouna, and Dieter Pfoser. Spatial auto-regressive dependency interpretable learning based on spatial topological constraints. ACM Transactions on Spatial Algorithms and Systems (TSAS), 5 (3):19:1–19:??, September 2019. CODEN ???? ISSN 2374-

0353. URL https://dl.acm.
org/citation.cfm?id=3339823.

Zhang:2021:ADI

[ZSBA21] Xiaowei Zhang, Aly Shehata, Bedrich Benes, and Daniel Aliaga. Automatic deep inference of procedural cities from global-scale spatial data. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7 (2):6:1–6:28, February 2021. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3423422.

Zhang:2020:DPS

[ZSFB20] Rui Zhang, Kevin G. Stanley, Daniel Fuller, and Scott Bell. Differentiating population spatial behavior using representative features of geospatial mobility (Re-FGeM). ACM Transactions on Spatial Algorithms and Systems (TSAS), 6(1):2:1–2:25, February 2020. CODEN ???? ISSN 2374-0353 (print), 2374-0361 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/abs/10.1145/3362063.

Zhou:2020:SST

[ZWT^{+}20] Fan Zhou, Hantao Wu, Goce Trajcevski, Ashfaq Khokhar, and Kunpeng Zhang. Semi-supervised trajectory understanding with POI attention for end-to-end trip recommendation. ACM Transactions on Spatial Algorithms and Systems (TSAS), 6(2):13:1– 13:25, February 2020. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

                                        URL https://dl.acm.org/doi/
                                        10.1145/3378890.

Zhu:2021:PCK

[ZYW^{+}21] Huaijie Zhu, Xiaochun Yang, Bin Wang, Wang-Chien Lee, Jian Yin, and Jianliang Xu. Processing continuous k nearest neighbor queries in obstructed space with Voronoi diagrams. ACM Transactions on Spatial Algorithms and Systems (TSAS), 7 (2):8:1–8:27, February 2021. CO-DEN ???? ISSN 2374-0353 (print), 2374-0361 (electronic).

URL https://dl.acm.org/doi/
10.1145/3425955.